A chemist studying the pollutant SO_2 (g) places 1
Answer #1

a) reaction is

2SO2 + O2   \rightarrow 2SO3

According to reaction 2 mole of SO2 react with 1 mole of O2 then for complete reaction 1 mole of SO2 require 0.5 mole of O2 thus, 0.5 mole of O2 remain unreacted

and 2 mole of SO2 produce 2 mole of SO3 then 1 mole of SO2 produce 1 mole of SO3

total mole of container after complete reaction = 0.5 mole O2 + 1 mole SO3 = 1.5 mole

to calculate final volume use ideal gas equation .

We know that PV = nRT

V = nRT/P

n = 1.5 mole,

T = 5090C = 509+273.15 = 782.15K,

P= 1.69 atm,

R = 0.08205 L atm mol-1 K-1 ( R = gas constant)

V = ?

Substitute these value in above equation.

V = 1.5 \times 0.08205 \times 782.15/1.69 = 56.96 L

final volume of gas is 56.96 L

b) calculate initial volume

initial mole = 1 mole SO2 + 1 mole O2 = 2 mole

to calculate initial volume use ideal gas equation .

We know that PV = nRT

V = nRT/P

n = 2 mole,

T = 5090C = 509+273.15 = 782.15K,

P= 1.69 atm,

R = 0.08205 L atm mol-1 K-1 ( R = gas constant)

V = ?

Substitute these value in above equation.

V = 2 \times  0.08205  \times 782.15/1.69 = 75.94 L

initial volume of gas is 75.94 L

nitial volume of gas is 75.94 L and final volume of gas is 56.96 L thus compession take place

therefore work done on the system.